OpenStudy (anonymous):

given f(x)=|2x-8|, find the value of f(1)-f(2). a.10 b.6 c.1 d.2

OpenStudy (mathstudent55):

To find f(1), replace x with 1 in the expression. f(1) = |2(1) - 8| Can you evaluate the right side?

OpenStudy (anonymous):

b?

OpenStudy (anonymous):

d

OpenStudy (mathstudent55):

I don't know. I'm not guessing an answer. Since you are asking about this problem, I'm trying to help you by breaking it down into steps. I asked you about the first step. How about answering my question and trying to solve it instead of guessing?

OpenStudy (mathstudent55):

Oh, I see. Perhaps your first answer was an answer to my question.

OpenStudy (anonymous):

I'm not guessing. I solved the first step but than solved the second and from there my answer is D.

OpenStudy (mathstudent55):

f(1) = |2(1) - 8| f(1) = |2 - 8| f(1) = |-6| = 6

OpenStudy (mathstudent55):

Yes, I think I figured that out.

OpenStudy (mathstudent55):

f(2) = |2(2) - 8| f(2) = |4 - 8| f(2) = |-4| = 4 So f(1) - f(2) = 6 - 4 = 2 You are correct. Sorry about my misunderstanding above.

OpenStudy (anonymous):

No sorry on my part and thank you!! I wish my book would break questions down the same way!

OpenStudy (mathstudent55):

If you have any more questions, keep asking questions and following the solutions. They may help you with this. After a while, you'll be able to break down the problem into simple steps.